# Sparse–Smooth Spatially Varying Coefficient Quantile Regression

### v1

@@ Appendix A Proofs of Theorems
Write $\Delta\bm{\alpha}=\hat{\bm{\alpha}}-\bm{\alpha}_{0}$, $\Delta\bm{\beta}_{G}=\hat{\bm{\beta}}_{G}-\bm{\beta}_{G0}$ and $\Delta\bm{\delta}_{j}=\hat{\bm{\delta}}_{j}-\bm{\delta}_{j0}$, and define the local perturbation $\bm{h}=\sum_{j=1}^{p}\bm{X}_{\odot j}\Delta\bm{\delta}_{j}$. By feasibility, each $\Delta\bm{\delta}_{j}\in\mathcal{S}$ (componentwise degree–weighted centering). Let $\bm{r}_{0}=\bm{\varepsilon}$ be the true residual vector. Using (15) with $t_{i}=h_{i}+\bm{Z}_{i}^{\top}\Delta\bm{\alpha}+\bm{X}_{i}^{\top}\Delta\bm{\beta}_{G}$ and summing, we obtain
$\sum_{i=1}^{n}\Big\{\rho_{\tau}(\varepsilon_{i}-t_{i})-\rho_{\tau}(\varepsilon_{i})\Big\}=-\langle\bm{\psi}_{\tau}(\bm{\varepsilon}),\bm{t}\rangle+\sum_{i=1}^{n}\int_{0}^{t_{i}}\big\{\mathbf{1}(\varepsilon_{i}\leq s)-\mathbf{1}(\varepsilon_{i}\leq 0)\big\}ds.$ (19)
By Assumption 7 (local quadratic expansion) and bounded covariates (Assumption 2), uniformly on events where $\|\Delta\bm{\theta}_{P}\|_{2}=O_{p}(n^{-1/2})$,
$\sum_{i=1}^{n}\Big\{\rho_{\tau}(\varepsilon_{i}-t_{i})-\rho_{\tau}(\varepsilon_{i})\Big\}=\tfrac{1}{2}\bm{t}^{\top}\bm{W}\bm{t}-\langle\bm{\psi}_{\tau}(\bm{\varepsilon}),\bm{t}\rangle+o_{p}(\|\bm{t}\|_{2}^{2}),$ (20)
where $\bm{W}=\operatorname{diag}\{f_{\varepsilon\mid Z,X,U}(0),\dots\}$ has entries bounded between $c_{f}$ and $C_{f}$.
Part (i): mean-squared rate.
Insert $(\bm{\theta}_{P},\bm{\Theta}_{L})=(\bm{\theta}_{P0},\bm{\Theta}_{L0})$ into (18). Combine (20) with polarization for the Laplacian term to obtain for every $j$,
$\tfrac{1}{2}\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{W}^{2}+\lambda_{2n}\Big(\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}-\hat{\bm{\delta}}_{j}^{\top}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}\Big)+\lambda_{1n}w_{j,n}\big(\|\bm{\delta}_{j0}\|_{2}-\|\hat{\bm{\delta}}_{j}\|_{2}\big)$
$\qquad\leq\langle\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon}),\Delta\bm{\delta}_{j}\rangle+R_{n,j},$ (21)
with $R_{n,j}=o_{p}(\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2})$ collecting the parametric perturbations and higher-order terms. Using
$\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}-\hat{\bm{\delta}}_{j}^{\top}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}=-\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\Delta\bm{\delta}_{j}-2\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0},$
the centering $\Delta\bm{\delta}_{j}\in\mathcal{S}$ and (17) yield
$\tfrac{1}{2}c_{f}\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}+\lambda_{2n}c_{\lambda}\|\Delta\bm{\delta}_{j}\|_{2}^{2}\ \leq\ \langle\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon}),\Delta\bm{\delta}_{j}\rangle+2\lambda_{2n}\big|\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big|+o_{p}(\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}).$ (22)
By Cauchy–Schwarz and Lemma-type bounds for $\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon})$ (bounded covariates, Assumption 2), the stochastic term is $O_{p}(\sqrt{n}\|\Delta\bm{\delta}_{j}\|_{2})$. The bias term satisfies
$\big|\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big|\leq\big(\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\Delta\bm{\delta}_{j}\big)^{1/2}\big(\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big)^{1/2}\leq C\|\Delta\bm{\delta}_{j}\|_{2}\mathcal{S}_{j}^{1/2},$
with $\mathcal{S}_{j}=\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}$ finite by Assumption 4. Absorbing the $o_{p}(\cdot)$ and dividing by $\|\Delta\bm{\delta}_{j}\|_{2}$,
$\lambda_{2n}c_{\lambda}\|\Delta\bm{\delta}_{j}\|_{2}\ \leq\ C\Big(\sqrt{n}+\lambda_{2n}\mathcal{S}_{j}^{1/2}\Big),$
so that
Finally, using boundedness of $X_{ij}$ (Assumption 2) to relate $\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}$ and $\|\Delta\bm{\delta}_{j}\|_{2}^{2}$, we obtain
and the bound is balanced by $\lambda_{2n}\asymp(n\mathcal{S}_{j})^{-1/2}$, proving (i).
Part (ii): group selection consistency.
Consider the KKT conditions for each $\hat{\bm{\delta}}_{j}$:
$\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})+2\lambda_{2n}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}+\bm{D}\hat{\bm{\mu}}_{j}\in-\lambda_{1n}w_{j,n}\partial\|\hat{\bm{\delta}}_{j}\|_{2},\qquad\bm{1}_{\mathcal{C}}^{\top}\bm{D}\hat{\bm{\delta}}_{j}=0.$
Inactive groups. Let $j\in\mathcal{A}_{0}$. Test $\hat{\bm{\delta}}_{j}=\bm{0}$. Projecting onto the centered subspace (which eliminates $\bm{D}\hat{\bm{\mu}}_{j}$) and using $\|\mathsf{Proj}_{D}(\bm{v})\|_{2}\leq\|\bm{v}\|_{2}$ yields
$\Big\|\mathsf{Proj}_{D}\big(\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\big)\Big\|_{2}\ \leq\ \|\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\|_{2}=O_{p}(\sqrt{n})\ \leq\ \lambda_{1n}w_{j,n}$
with probability tending to one because $\sqrt{n}\lambda_{1n}\to\infty$ and $\inf_{j\in\mathcal{A}_{0}}w_{j,n}\geq c_{w}>0$ by Assumption 5. Hence $\Pr(\hat{\bm{\delta}}_{j}=\bm{0})\to 1$ for each $j\in\mathcal{A}_{0}$.
Active groups. Let $j\in\mathcal{A}_{0}^{c}$. Consider the restricted problem that sets $\bm{\delta}_{j}=\bm{0}$ for $j\in\mathcal{A}_{0}$ and optimizes over $\{\bm{\delta}_{j}:j\in\mathcal{A}_{0}^{c}\}$ and $\bm{\theta}_{P}$. Under Assumptions 56, the groupwise beta–min separation and bounded adaptive weights on $\mathcal{A}_{0}^{c}$ ensure that the restricted solution satisfies $\|\hat{\bm{\delta}}_{j}^{\mathrm{res}}\|_{2}\geq c_{\beta}\lambda_{1n}w_{j,n}-o_{p}(\lambda_{1n})>0$ for $j\in\mathcal{A}_{0}^{c}$, while the subgradient conditions for $j\in\mathcal{A}_{0}$ hold with strict inequality (dual feasibility). Therefore the restricted solution is also optimal for the full problem (primal–dual witness, cf. [YuanLin2006, Simon2013SGL]), implying $\hat{\bm{\delta}}_{j}\neq\bm{0}$ for all $j\in\mathcal{A}_{0}^{c}$. On this event, the rate in part (i) holds for $j\in\mathcal{A}_{0}^{c}$. ∎
Proof of Lemma 1.
Let $e_{j,i}=\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})$ and write $\bm{e}_{j}=\hat{\bm{\delta}}_{j}-\bm{\delta}_{j0}$. Using Cauchy–Schwarz, boundedness of $f_{\varepsilon\mid Z,X,U}(0)$ and of covariates (Assumption 2), and denoting $\bm{G}_{i}=(\bm{Z}_{i}^{\top},\bm{X}_{i}^{\top})^{\top}$,
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}\right\|_{2}$ $\leq C\left(\frac{1}{n}\sum_{i=1}^{n}\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)^{2}\right)^{1/2}\left(\frac{1}{n}\sum_{i=1}^{n}\|\bm{G}_{i}\|_{2}^{2}\right)^{1/2}$
$\leq C^{\prime}\left(\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}e_{j,i}^{2}\right)^{1/2}=C^{\prime}\left(\frac{1}{p}\sum_{j=1}^{p}\frac{1}{n}\sum_{i=1}^{n}e_{j,i}^{2}\right)^{1/2}.$
By Theorem 3(i), for each $j$,
$\frac{1}{n}\sum_{i=1}^{n}e_{j,i}^{2}=O_{p}\Big(\frac{1}{n\lambda_{2n}}+\lambda_{2n}\mathcal{S}_{j}\Big).$
Averaging over $j$ preserves the order (with $\bar{\mathcal{S}}=p^{-1}\sum_{j}\mathcal{S}_{j}$), hence
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}\right\|_{2}=O_{p}\left(\Big(\tfrac{1}{n\lambda_{2n}}+\lambda_{2n}\bar{\mathcal{S}}\Big)^{1/2}\right).$
For the canonical balancing $\lambda_{2n}\asymp n^{-1/2}$, the right-hand side is $O_{p}(n^{-1/4})=o_{p}(1)$, which proves the claim. ∎
Proof of Theorem 4.
@@
Let $\hat{\bm{r}}$ be the fitted residuals and $r_{0,i}=\varepsilon_{i}$. The stationarity of the parametric block gives
$\bm{0}=\sum_{i=1}^{n}\psi_{\tau}(\hat{r}_{i})\bm{G}_{i}=\sum_{i=1}^{n}\psi_{\tau}(r_{0,i})\bm{G}_{i}+\sum_{i=1}^{n}\big\{\psi_{\tau}(\hat{r}_{i})-\psi_{\tau}(r_{0,i})\big\}\bm{G}_{i}.$
Using the linearization
$\psi_{\tau}(\varepsilon_{i}-t_{i})-\psi_{\tau}(\varepsilon_{i})=-f_{\varepsilon\mid Z,X,U}(0)t_{i}+\eta_{i}(t_{i}),$
with $\sum_{i}\eta_{i}(t_{i})\bm{G}_{i}=o_{p}(\sqrt{n}\|\bm{t}\|_{2})$ uniformly for $\|\bm{t}\|_{2}=O_{p}(1)$ (see [Koenker2005, Sec. 4.3]), we obtain
$\bm{0}$ $=\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}-\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)t_{i}\bm{G}_{i}+o_{p}(\sqrt{n}),$ (23)
$t_{i}$ $=\bm{Z}_{i}^{\top}\Delta\bm{\alpha}+\bm{X}_{i}^{\top}\Delta\bm{\beta}_{G}+\sum_{j=1}^{p}X_{ij}\big\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\big\}.$ (24)
Divide (23) by $\sqrt{n}$ and rearrange the terms that involve $(\Delta\bm{\alpha},\Delta\bm{\beta}_{G})$:
$\sqrt{n}\bm{M}\left[\begin{array}[]{c}\Delta\bm{\alpha}\\ \Delta\bm{\beta}_{G}\end{array}\right]=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}+o_{p}(1),$
where $\bm{M}=\mathrm{E}[f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}]$ (Assumption 7) and $e_{j,i}=\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})$. By Lemma 1, the middle term on the right is $o_{p}(1)$ after multiplication by $\bm{M}^{-1}$. Therefore
$\sqrt{n}\left[\begin{array}[]{c}\Delta\bm{\alpha}\\ \Delta\bm{\beta}_{G}\end{array}\right]=\bm{M}^{-1}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}\right)+o_{p}(1).$
By a multivariate central limit theorem, the leading sum converges in distribution to $\mathcal{N}(\bm{0},\bm{V})$ with $\bm{V}=\tau(1-\tau)\mathrm{E}(\bm{G}_{i}\bm{G}_{i}^{\top})$, yielding
$\sqrt{n}(\hat{\bm{\theta}}_{P}-\bm{\theta}_{P0})\ \overset{d}{\longrightarrow}\ \mathcal{N}(\bm{0},\ \bm{M}^{-1}\bm{V}\bm{M}^{-1}).$
Oracle property. By Theorem 3(ii), $\Pr(\hat{\mathcal{A}}=\mathcal{A}_{0})\to 1$. On this event, $\hat{\bm{\delta}}_{j}=\bm{0}$ for $j\in\mathcal{A}_{0}$ and the restricted problem for $(\bm{\theta}_{P},\{\bm{\delta}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ coincides with the oracle problem that knows $\mathcal{A}_{0}$. Hence the joint limit of $(\hat{\bm{\theta}}_{P},\{\hat{\bm{\delta}}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ equals the oracle limit distribution. ∎

### v2

@@ Appendix A Proofs of Theorems
We first derive the asymptotic distribution of the parametric block $\hat{\bm{\theta}}_{P}=(\hat{\bm{\alpha}}^{\top},\hat{\bm{\beta}}_{G}^{\top})^{\top}$ and then argue the oracle property.
@@
Let $\hat{r}_{i}$ be the fitted residuals and recall $\varepsilon_{i}=r_{i}(\bm{\theta}_{P0},\bm{\Theta}_{L0})$. The KKT conditions for the parametric block read
$\sum_{i=1}^{n}\psi_{\tau}(\hat{r}_{i})\bm{G}_{i}=\bm{0}.$
Write $\Delta\bm{\theta}_{P}=\hat{\bm{\theta}}_{P}-\bm{\theta}_{P0}$. Using the representation
$\hat{r}_{i}=\varepsilon_{i}-\bm{G}_{i}^{\top}\Delta\bm{\theta}_{P}-\sum_{j=1}^{p}X_{ij}\left\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\right\},$
and applying the same linearization as in (A) (see also [Koenker2005, Sec. 4.3]), we obtain
$\psi_{\tau}(\hat{r}_{i})-\psi_{\tau}(\varepsilon_{i})=-f_{\varepsilon\mid Z,X,U}(0)\Big(\bm{G}_{i}^{\top}\Delta\bm{\theta}_{P}+\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)+\eta_{i},$ (27)
where the remainder $\eta_{i}$ satisfies $\sum_{i=1}^{n}\eta_{i}\bm{G}_{i}=o_{p}(\sqrt{n}\|\Delta\bm{\theta}_{P}\|_{2})+o_{p}(1)$ uniformly on events where $\|\Delta\bm{\theta}_{P}\|_{2}=O(n^{-1/2})$. Summing (27) over $i$ and rearranging terms yields
$\bm{0}$ $=\sum_{i=1}^{n}\psi_{\tau}(\hat{r}_{i})\bm{G}_{i}$
$=\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}-\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}\Delta\bm{\theta}_{P}-\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}$
$\hskip 284.52756pt+\sum_{i=1}^{n}\eta_{i}\bm{G}_{i}.$
Dividing by $\sqrt{n}$ and denoting
$\bm{M}_{n}=\frac{1}{n}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top},\quad\bm{S}_{n}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i},$
we can rewrite this as
$\sqrt{n}\bm{M}_{n}\Delta\bm{\theta}_{P}=\bm{S}_{n}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}+o_{p}(1).$
Assumption 7 states that $\bm{M}_{n}\to\bm{M}$ in probability and that $\bm{M}$ is positive definite. By Lemma 1, the second term on the right-hand side converges to zero in probability. Thus
$\sqrt{n}\Delta\bm{\theta}_{P}=\bm{M}^{-1}\bm{S}_{n}+o_{p}(1).$
By the multivariate central limit theorem and Assumption 7,
$\bm{S}_{n}\ \overset{d}{\longrightarrow}\ \mathcal{N}(\bm{0},\bm{V}),\quad\bm{V}=\tau(1-\tau)\mathrm{E}(\bm{G}_{i}\bm{G}_{i}^{\top}),$
and Slutsky’s theorem gives
$\sqrt{n}(\hat{\bm{\theta}}_{P}-\bm{\theta}_{P0})\overset{d}{\longrightarrow}\mathcal{N}\big(\bm{0},\bm{M}^{-1}\bm{V}\bm{M}^{-1}\big),$
which establishes the first part of the theorem.
We now argue the oracle property. Let $\hat{\mathcal{A}}$ denote the estimated set of active (truly varying) covariates, i.e. $\hat{\mathcal{A}}=\{j:\hat{\bm{\delta}}_{j}\neq\bm{0}\}$. By Theorem 3(ii),
$\Pr(\hat{\mathcal{A}}=\mathcal{A}_{0})\to 1.$
On the event $\{\hat{\mathcal{A}}=\mathcal{A}_{0}\}$ the minimization problem defining $(\hat{\bm{\theta}}_{P},\hat{\bm{\Theta}}_{L})$ reduces exactly to the oracle problem in which we a priori set $\bm{\delta}_{j}\equiv\bm{0}$ for all $j\in\mathcal{A}_{0}$ and optimize over $(\bm{\theta}_{P},\{\bm{\delta}_{j}:j\in\mathcal{A}_{0}^{c}\})$. Therefore the joint limit distribution of $(\hat{\bm{\theta}}_{P},\{\hat{\bm{\delta}}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ coincides with that of the corresponding oracle estimator. In particular, the parametric block enjoys the same asymptotic linear representation and covariance as if the true active set were known. ∎
Finally, we briefly comment on variance estimation for the limiting distribution. A plug-in estimator of the asymptotic covariance matrix $\bm{M}^{-1}\bm{V}\bm{M}^{-1}$ is obtained by replacing $\bm{M}$ and $\bm{V}$ with
$\hat{\bm{M}}=\frac{1}{n}\sum_{i=1}^{n}\hat{f}_{i}(0)\bm{G}_{i}\bm{G}_{i}^{\top},\quad\hat{\bm{V}}=\tau(1-\tau)\frac{1}{n}\sum_{i=1}^{n}\bm{G}_{i}\bm{G}_{i}^{\top},$
where $\hat{f}_{i}(0)$ is either a kernel estimate of the density of $\varepsilon_{i}$ at zero based on residuals, or an implicit estimate based on a smoothed check loss with bandwidth $h_{n}\downarrow 0$ and $\sqrt{n}h_{n}\to\infty$. If spatial dependence among the observations is of concern, one can further use spatial block bootstrap or heteroskedasticity-and-autocorrelation consistent (HAC) corrections in the sandwich step, but this lies beyond the main focus of the present paper.